# Clustering by Mining Density Distributions and Splitting Manifold Structure

### v1

@@ Abstract
Spectral clustering requires the time-consuming decomposition of the Laplacian matrix of the similarity graph, thus limiting its applicability to large datasets. To improve the efficiency of spectral clustering, a top-down approach was recently proposed, which first divides the data into several micro-clusters (granular-balls), then splits these micro-clusters when they are not “compact”, and finally uses these micro-clusters as nodes to construct a similarity graph for more efficient spectral clustering. However, this top-down approach is challenging to adapt to unevenly distributed or structurally complex data. This is because constructing micro-clusters as a rough ball struggles to capture the shape and structure of data in a local range, and the simplistic splitting rule that solely targets “compactness” is susceptible to noise and variations in data density and leads to micro-clusters with varying shapes, making it challenging to accurately measure the similarity between them. To resolve these issues, this paper first proposes to start from local structures to obtain micro-clusters, such that the complex structural information inside local neighborhoods is well captured by them. Moreover, by noting that Euclidean distance is more suitable for convex sets, this paper further proposes a data splitting rule that couples local density and data manifold structures, so that the similarities of the obtained micro-clusters can be easily characterized. A novel similarity measure between micro-clusters is then proposed for the final spectral clustering. A series of experiments based on synthetic and real-world datasets demonstrate that the proposed method has better adaptability to structurally complex data than granular-ball based methods.
@@ Introduction
Spectral Clustering is a representative graph partition clustering algorithm, the core idea of which is to (approximately) minimize a cut loss of partitioning a similarity graph of data by removing some edges. It has attracted wide attention due to its exceptionally good performance in handling non-convex shaped clusters (von Luxburg 2007; Chen et al. 2011; He et al. 2019).
@@
To address these issues, this paper proposes another approach to represent data in a more coarse granularity. It first tries to capture more detailed local structural information of data by constructing micro-clusters from an estimation of data densities. Then it follows a more sophisticated splitting rule that also considers the convexity of data (called manifold curvature) to improve the quality of micro-clusters. The more convex distribution of data in micro-clusters also makes the design of similarity measure more straightforward and thus the final Spectral Clustering more effective.
@@
Another direction is to use anchors to characterize data, where the similarity between points are approximated using the similarity between point and anchors. Since the similarity graph between point and anchors is a bipartite graph, the spectral decomposition can be more efficient by working on a smaller matrix. For example, Cai and Chen (2015) performed K-means to obtain anchor points and constructed a sparsified bipartite graph between data points and these anchors by keeping only the connections of several nearest anchors for each point. Huang et al. (2020) further improved computational efficiency by first applying K-means on randomly sampled points to obtain anchors, then constructing a bipartite graph via fast approximate nearest neighbors, and finally performing transfer-cut on the bipartite graph for efficient spectral decomposition. There are also works (Gao et al. 2024; Nie et al. 2024) on combining anchors with the optimization process to increase efficiency.
@@
Given a dataset $X=\{x_{1},...,x_{n}\}$ $(x_{i}\in\mathbb{R}^{d})$, the target of GBSC (Xie et al. 2023) is to use a set of granular-balls as micro-clusters to cover the dataset, such that each data point belongs to a single granular-ball. A granular-ball is just an $d$ dimensional ball while the radius of the balls can vary from each other.
Specifically, suppose $GB_{j}$ is a granular-ball covering the $m$ data points $\{x_{i_{1}},\ldots,x_{i_{m}}\}\subseteq X$, then the center $c_{j}$ and the radius of $GB_{j}$ are determined as $c_{j}=\frac{1}{m}\sum_{s=1}^{m}x_{i_{s}}$ and $r_{j}=\max_{s}(\begin{Vmatrix}x_{i_{s}}-c_{j}\end{Vmatrix})$, where |||||$\cdot$||| | ⋅ | | denotes the $\ell_{2}$ norm. Note that if the radius of the granular-ball is large then the “granularity” is coarse and the clustering on the balls would be fast while much structural information would be lost; otherwise, the “granularity” is fine but clustering would be slow. Therefore, generating granular-balls needs to balance granularity and the quality of the balls.
@@
The generation of granular-balls in a top-down manner work as follows. First, a granular-ball $GB_{A}$ covering the entire dataset is generated. Then, two farthest points in $p_{1}$ and $p_{2}$ are selected, and the points in $GB_{A}$ are assigned to $p_{1}$ if they are closer to $p_{1}$ than to $p_{2}$. Two children balls $GB_{A_{1}}$ and $GB_{A_{2}}$ are then generated using the two subsets of points.
@@
The density $\rho(x_{i})$ of a point $x_{i}$, is defined as
$\rho(x_{i})=\sum_{j=1}^{k}\exp(-dist_{ij}^{2}),$ (3)
where $dist_{ij}$ refers to the Euclidean distance between point $x_{i}$ and its neighbor $x_{j}$.
@@
where the set $\mathcal{H}(x_{i})=\{x_{j}\mid x_{j}\in N_{k}(x_{i}),\rho(x_{j})>\rho(x_{i})\}$. Points without a leader are called core points, and the set of core points is denoted as $\mathsf{core}=\{x_{i}\mid\mathsf{leader}(x_{i})=\text{None}\}$.
@@
Fig. 3 illustrates the pseudo-clusters of Jain with $k=10$ and Spiral with $k=4$. It can be seen that pseudo-clusters better reflects the local characteristics of the dataset, and the problem of a granular-ball in Fig. 1 containing points from different clusters has been rectified.
@@
On the left side of Fig. 3, the pseudo-clusters of numbers 0, 1, 2, 3, and 8 exhibit non-convex structures and complex shapes, and the distance between these pseudo-clusters and others can not be easily measured with Euclidean based distances. For example, for pseudo-cluster 2, although its core point is close to pseudo-cluster 1, the points in the left part of it is far away. The situation is similar on the other side.
@@
Intuitively, in a pseudo-cluster, the shortest path distance reflects the geodesic distance between points, while the Euclidean distance represents the “direct” distance between them. Thus, we define the manifold curvature as the ratio of the shortest path distance to the Euclidean distance between the endpoints. The larger the ratio is, the more “curved” the pseudo-cluster will be. When the ratio approaches 1, the data will be nearly convex.
@@
To determine whether a pseudo-cluster $p_{t}$ should be split, we consider both the manifold curvature and the compactness of the original pseudo-cluster and its child pseudo-clusters according to Eq. 1 and Eq. 2. In particular, if $MC(p_{t})\geq\lambda$ and $DM_{weight}<DM(p_{t})$, then $p_{t}$ will be split. In order to avoid too small pseudo-clusters, we also require $p_{t}$ should contain a minimum number $\beta$ of points, as in (Xie et al. 2023).
@@
0:  Dataset $X$, the number of nearest neighbors $k$, the manifold curvature threshold $\lambda$, and the mimum size of pseudo-cluster $\beta$.
@@
5:     for $x_{i},x_{j}\in p_{t}$ do
6:        Calculate $path\_dist_{ij}$ on $T(p_{t})$;
7:     end for
8:  end for
9:  while $\exists p_{t}$ s.t. $MC(p_{t})>\lambda$ and $DM_{weight}<DM(p_{t})$ and $|p_{t}|>\beta$ do
10:     Split $p_{t}$ into $p_{t_{1}}$ and $p_{t_{2}}$ and add them to $\mathsf{pseudo}\_\mathsf{clusters}$;
11:  end while
12:  Calculate similarity matrix $S$;
13:  Perform Spectral Clustering on the $S$ to obtain final clustering results.
@@
where $\mathsf{SNN}(p_{i},p_{j})=\cup_{x\in p_{i}}N_{k}(x)\cap\cup_{x\in p_{j}}N_{k}(x)$ is the shared nearest neighbors of $p_{i}$ and $p_{j}$, and $c\_dist(p_{i},p_{j})$ is the Euclidean distance between the centroids of $p_{i}$ and $p_{j}$. A centroid of $p_{i}$ is $\frac{1}{m_{i}}\sum_{x\in p_{i}}x$, which is not necessarily a core point.
@@
Suppose the dataset $X$ has $n$ samples with the dimensionality of $d$, the final number of pseudo-clusters is $m$, and the number of nearest neighbors is $k$. In Algorithm 1, the time complexity for finding k-nearest neighbors using the KD-tree method is $\mathcal{O}((k+n)\log n)$, the time complexity for calculating density is $\mathcal{O}(kn)$, and the time complexity for finding the leader point of a data point is $\mathcal{O}(kn)$.
@@
Actually, $\sum_{i}\mathcal{O}(n_{i}^{2}\log n_{i})$ is bounded by $\mathcal{O}(nn_{*}\log n_{*})$, where $n_{*}$ is the maximum of $n_{i}$, as shown in the following theorem (see supplementary material for proof). Therefore, the overall time complexity of the algorithm is thus $\mathcal{O}(nn_{*}(\log n_{*})^{2}+m^{3})$.
@@
We used three evaluation metrics: ARI (Steinley 2004), NMI (Xu, Liu, and Gong 2003), and ACC (Yang et al. 2010) for clustering analysis. Since the proposed algorithm is based on local density peaks and pseudo-cluster splitting followed by spectral clustering, we selected the following eight comparison algorithms: GBSC (Xie et al. 2023), GB-DP (Cheng et al. 2023), LDP-MST (Cheng et al. 2021), USPEC (Huang et al. 2020), SC (Shi and Malik 2000), DPC (Rodriguez and Laio 2014), DEMOS (Guan et al. 2023), and DPC-DBFN (Lotfi, Moradi, and Beigy 2020).
For MDMSC, we searched for $k$ in the range of 2 to 50 and for $\beta$ in $\{8,16\}$, and the threshold $\lambda$ was set to $1.5$. The implementation details of other algorithms are provided in the supplementary material.
@@
In terms of datasets, we used 4 synthetic datasets and 13 real-world datasets (Table 1). All datasets were normalized, and the results were averaged over 10 runs.
@@
Table 2 shows the results (bold means best) of our algorithm and the other eight comparison algorithms on real-world datasets. MDMSC achieves the best performance on most datasets. For example, on the high-dimensional dataset pengleukEW, MDMSC significantly outperforms most comparison algorithms. We attribute this to the consideration of manifold curvature, allowing it to effectively handle complex high-dimensional datasets. Additionally, on the large-scale dataset Pendigits, our algorithm also shows superior performance, demonstrating its capability to effectively handle large-scale datasets by partitioning pseudo-clusters. Although our algorithm does not always achieve the best results on a few datasets like landsatEW and soybean_test, its performance is very close to the optimal ones. This indicates that our algorithm possesses strong adaptability and effectiveness across a wide range of application scenarios.
We also performed Friedman test and subsequent Nemenyi test on the ACC results. Friedman test showed that there exists significant differences ($p<0.05$) between the results and Nemenyi test confirmed that our algorithm significantly outperforms all of the other algorithms: $CD=2.10$ and the rank differences of MDMSC with other algorithms are all larger than $2.10$.

### v2

@@ Abstract
Spectral clustering requires the time-consuming decomposition of the Laplacian matrix of the similarity graph, thus limiting its applicability to large datasets. To improve the efficiency of spectral clustering, a top-down approach was recently proposed, which first divides the data into several micro-clusters (granular-balls), then splits these micro-clusters when they are not “compact”, and finally uses these micro-clusters as nodes to construct a similarity graph for more efficient spectral clustering. However, this top-down approach is challenging to adapt to unevenly distributed or structurally complex data. This is because constructing micro-clusters as a rough ball struggles to capture the shape and structure of data in a local range, and the simplistic splitting rule that solely targets “compactness” is susceptible to noise and variations in data density and leads to micro-clusters with varying shapes, making it challenging to accurately measure the similarity between them. To resolve these issues and improve spectral clustering, this paper first proposes to start from local structures to obtain micro-clusters, such that the complex structural information inside local neighborhoods is well captured by them. Moreover, by noting that Euclidean distance is more suitable for convex sets, this paper further proposes a data splitting rule that couples local density and data manifold structures, so that the similarities of the obtained micro-clusters can be easily characterized. A novel similarity measure between micro-clusters is then proposed for the final spectral clustering. A series of experiments based on synthetic and real-world datasets demonstrate that the proposed method has better adaptability to structurally complex data than granular-ball based methods.
@@ Introduction
Spectral lustering is a representative graph partition clustering algorithm, the core idea of which is to (approximately) minimize a cut loss of partitioning a similarity graph of data by removing some edges. It has attracted wide attention due to its exceptionally good performance in handling non-convex shaped clusters (von Luxburg 2007; Chen et al. 2011; He et al. 2019).
@@
To resolve these issues and accelerate spectral clustering, this paper proposes another approach to represent data in a more coarse granularity. It first tries to capture more detailed local structural information of data by constructing micro-clusters from an estimation of data densities. Then it follows a more sophisticated splitting rule that also considers the convexity of data (called manifold curvature) to improve the quality of micro-clusters. The more convex distribution of data in micro-clusters also makes the design of similarity measure more straightforward and thus the final spectral clustering more effective.
@@
Another direction is to use anchors to characterize data, where the similarity between points are approximated using the similarity between points and anchors. Since the similarity graph between points and anchors is a bipartite graph, the spectral decomposition can be more efficient by working on a smaller matrix. For example, Cai and Chen (2015) performed K-means to obtain anchor points and constructed a sparsified bipartite graph between data points and these anchors by keeping only the connections of several nearest anchors for each point. Huang et al. (2020) further improved computational efficiency by first applying K-means on randomly sampled points to obtain anchors, then constructing a bipartite graph via fast approximate nearest neighbors, and finally performing transfer-cut on the bipartite graph for efficient spectral decomposition. There are also works (Gao et al. 2024; Nie et al. 2024) on combining anchors with the optimization process to increase efficiency.
@@
Given a dataset $X=\{x_{1},...,x_{n}\}$ $(x_{i}\in\mathbb{R}^{d})$, the target of GBSC (Xie et al. 2023) is to use a set of granular-balls as micro-clusters to cover the dataset, such that each data point belongs to a single granular-ball. A granular-ball is just an $d$ dimensional ball while the radius of the ball can vary from each other.
Specifically, suppose $GB_{j}$ is a granular-ball covering the $m$ data points $\{x_{j_{1}},\ldots,x_{j_{m}}\}\subseteq X$, then the center $c_{j}$ and the radius of $GB_{j}$ are determined as $c_{j}=\frac{1}{m}\sum_{s=1}^{m}x_{j_{s}}$ and $r_{j}=\max_{s}(\begin{Vmatrix}x_{j_{s}}-c_{j}\end{Vmatrix})$, where |||||$\cdot$||| | ⋅ | | denotes the $\ell_{2}$ norm. Note that if the radius of the granular-ball is large then the “granularity” is coarse and the clustering on the balls would be fast while much structural information would be lost; otherwise, the “granularity” is fine but clustering would be slow. Therefore, generating granular-balls needs to balance granularity and the quality of the balls.
@@
The generation of granular-balls in a top-down manner works as follows. First, a granular-ball $GB_{A}$ covering the entire dataset is generated. Then, two farthest points $p_{1}$ and $p_{2}$ are selected, and the points in $GB_{A}$ are assigned to $p_{1}$ if they are closer to $p_{1}$ than to $p_{2}$. Two children balls $GB_{A_{1}}$ and $GB_{A_{2}}$ are then generated using the two subsets of points.
@@
The density $\rho_{k}(x_{i})$ of a point $x_{i}$, is defined as
$\rho_{k}(x_{i})=\sum_{j=1}^{k}\exp(-dist_{ij}^{2}),$ (3)
where $dist_{ij}$ refers to the Euclidean distance between point $x_{i}$ and its neighbor $x_{j}$. The density is measured by the sum of Gaussian kernels of the Euclidean distances, reflecting the compactness of the local structure around the point.
@@
where the set $\mathcal{H}(x_{i})=\{x_{j}\mid x_{j}\in N_{k}(x_{i}),\rho_{k}(x_{j})>\rho_{k}(x_{i})\}$. Points without a leader are called core points, and the set of core points is denoted as $\mathsf{core}=\{x_{i}\mid\mathsf{leader}(x_{i})=\text{None}\}$.
@@
Fig. 3 illustrates the pseudo-clusters of Jain with $k=10$ and Spiral with $k=4$. It can be seen that pseudo-clusters better reflect the local characteristics of the dataset, and the problem of a granular-ball in Fig. 1 containing points from different clusters has been rectified.
@@
On the left side of Fig. 3, the pseudo-clusters of numbers 1, 2, 3, and 4 exhibit non-convex structures and complex shapes, and the distance between these pseudo-clusters and others can not be easily measured with Euclidean based distances. For example, for pseudo-cluster 2, although its core point is close to pseudo-cluster 1, the points in the left part of it are far away. The situation is similar on the other side.
@@
Intuitively, if the geodesic distance (the shortest distance on a manifold) equals the Euclidean distance, then the set is convex, and non-convex otherwise. Here, we use the shortest path distance in MST to approximate the geodesic distance, and the ratio of them to measure curvature.
Thus, we define the manifold curvature as the ratio of the shortest path distance to the Euclidean distance between the endpoints. The larger the ratio is, the more “curved” the pseudo-cluster will be. When the ratio approaches 1, the data will be nearly convex.
@@
To determine whether a pseudo-cluster $p_{t}$ should be split, we consider both the manifold curvature and the compactness of the original pseudo-cluster and its child pseudo-clusters according to Eq. 1 and Eq. 2. In particular, if $MC(p_{t})\geq\lambda$ and $DM_{weight}<DM(p_{t})$, then $p_{t}$ will be split. Because geodesic distance is estimated by graph distance, the two distances would be equal only with sufficient sampling; otherwise, graph distance is larger than the true geodesic distance. Thus, we set a threshold of 1.5 for the $\lambda$. In order to avoid too small pseudo-clusters, we also require $p_{t}$ should contain a minimum number $\beta$ of points, as in (Xie et al. 2023).
@@
0:  Dataset $X$, the number of nearest neighbors $k$, the manifold curvature threshold $\lambda$, and the minimum size of pseudo-cluster $\beta$.
@@
5:     Calculate $DM_{weight}$, $DM(p_{t})$ and $MC(p_{t})$;
6:  end for
7:  while $\exists p_{t}$ s.t. $MC(p_{t})>\lambda$ and $DM_{weight}<DM(p_{t})$ and $|p_{t}|>\beta$ do
8:     Split $p_{t}$ into $p_{t_{1}}$ and $p_{t_{2}}$ and add them to $\mathsf{pseudo}\_\mathsf{clusters}$;
9:     Remove $p_{t}$ from $\mathsf{pseudo}\_\mathsf{clusters}$;
10:  end while
11:  Calculate similarity matrix $S$;
12:  Perform spectral clustering on the $S$ to obtain final clustering results.
@@
where $\mathsf{SNN}(p_{i},p_{j})=(\cup_{x\in p_{i}}N_{k}(x))\cap(\cup_{x\in p_{j}}N_{k}(x)$) is the shared nearest neighbors of $p_{i}$ and $p_{j}$, and $c\_dist(p_{i},p_{j})$ is the Euclidean distance between the centroids of $p_{i}$ and $p_{j}$. A centroid of $p_{i}$ is $\frac{1}{m_{i}}\sum_{x\in p_{i}}x$, which is not necessarily a core point.
@@
Suppose the dataset $X$ has $n$ samples with the dimensionality of $d$, the final number of pseudo-clusters is $m$, and the number of nearest neighbors is $k$. In Algorithm 1, the time complexity for finding k-nearest neighbors using the KD-tree method is $\mathcal{O}((d+k)n\log n)$, the time complexity for calculating density is $\mathcal{O}(kn)$, and the time complexity for finding the leader points is $\mathcal{O}(kn)$.
@@
Actually, $\sum_{i}\mathcal{O}(n_{i}^{2}\log n_{i})$ is bounded by $\mathcal{O}(nn_{*}\log n_{*})$, where $n_{*}$ is the maximum of $n_{i}$, as shown in the following theorem (see the proof). Therefore, the overall time complexity of the Algorithm 2 is thus $\mathcal{O}(nn_{*}(\log n_{*})+m^{3})$.
@@
Proof.
Note that $a_{1}^{2}\log a_{1}+a_{2}^{2}\log a_{2}\leq(a_{1}+a_{2})^{2}\log(a_{1}+a_{2})$ and more generally $a_{1}^{2}\log a_{1}+\ldots+a_{s}^{2}\log a_{s}\leq(a_{1}+\ldots a_{s})^{2}\log(a_{1}+\ldots+a_{s})$ for any $a_{i}\geq 1$.
Therefore, we can group the pseudo-clusters into $C$ groups such that the total number of points in each group is in $[n_{*}/2,n_{*}]$, with the possible exception of one group whose total number of points is less than $n_{*}/2$. Then there are at most $12\lceil n/n_{*}\rceil+$ 1 groups.
For each group containing the pseudo-clusters $p_{i_{1}},\ldots,p_{i_{s}}$, we have that the corresponding summation terms $n_{i_{1}}^{2}\log{n_{i_{1}}}+\ldots+n_{i_{s}}^{2}\log{n_{i_{s}}}\leq(n_{i_{1}}+\ldots+n_{i_{s}})^{2}\log{(n_{i_{1}}+\ldots+n_{i_{s}})}$. Because of the grouping strategy, $n_{i_{1}}+\ldots+n_{i_{s}}\leq n_{*}$, the summation of these terms is bounded by $n_{*}^{2}\log n_{*}$ and $\sum_{i}n_{i}^{2}\log n_{i}\leq Cn_{*}^{2}\log n_{*}$. Since $C\leq 2\lceil n/n_{*}\rceil+1$, $\sum_{i}n_{i}^{2}\log n_{i}\leq(2\lceil n/n_{*}\rceil+1)n_{*}^{2}\log{n_{*}}=\mathcal{O}(nn_{*}\log{n_{*}})$.
@@
The experimental environment is: Windows 10, CPU i5-9300H, GPU NVIDIA GeForce GTX 1050 and 8GB RAM, Python 3.9.
The experiments are conducted on 4 synthetic datasets: Aggregation (Gionis, Mannila, and Tsaparas 2007), Spiral (Chang and Yeung 2008), Jain (Jain and Law 2005), db2 (Ester et al. 1996) and 13 UCI datasets  [^1] https://archive.ics.uci.edu/. We used three evaluation metrics: ARI (Steinley 2004), NMI (Xu, Liu, and Gong 2003), and ACC (Yang et al. 2010) for clustering analysis. Since the proposed algorithm is based on local density peaks and pseudo-cluster splitting followed by spectral clustering, we selected the following eight comparison algorithms: GBSC (Xie et al. 2023), GB-DP (Cheng et al. 2023), LDP-MST (Cheng et al. 2021), USPEC (Huang et al. 2020), spectral clustering (SC) (Shi and Malik 2000), DPC (Rodriguez and Laio 2014), DEMOS (Guan et al. 2023), and DPC-DBFN (Lotfi, Moradi, and Beigy 2020).
The implementations of GBSC, GB-DP, LDP-MST, USPEC, DEMOS, and DPC-DBFN are from the source code provided by the authors. The implementation of SC is from the scikit-learn library, and the implementation of DPC is based on the algorithm described in the paper. For a fair comparison, we performed a search for the optimal hyperparameters required by each algorithm. Among them, GBSC and GB-DP do not require hyperparameter settings. U-SPEC has a hyperparameter $p$; if the number of data points $n$ is greater than 1000, we set $p$ to 1,000 and to $n$ ($n$ represents the number of samples in the data) otherwise. For LDP-MST, we followed the experimental setup from the original paper and applied PCA to reduce the dimensionality of datasets with more than 10 dimensions, selecting the best result between 2 and 10 dimensions. For SC, we used the library function to construct similarity using k-nearest neighbors. For DPC, the suggested average number of points within the cutoff distance should be 1% to 2% of the total data. We searched this ratio from 1.0 to 2.0 with a step size of 0.1 to find the optimal result. For DEMOS, we manually drew rectangular boxes in the decision graph to select cluster center points.The value of $k$ for the number of nearest neighbors is set according to $\sqrt{n}$ ($n$ represents the number of samples in the data) as described in the original paper. For DPC-DBFN, we searched for the optimal $k$ in the range of 2 to 100. For MDMSC, we searched for $k$ in the range of 2 to 50 and for $\beta$ in $\{8,16\}$, and the threshold $\lambda$ was set to $1.5$.
All datasets in Table 1 were normalized, and the results were averaged over 10 runs. The implementation of MDMSC can be found at https://github.com/SWJTU-ML/MDMSC.
@@
DataSet Ours GBSC GB-DP LDP-MST USPEC SC DPC DEMOS DPC-DBFN
Par. $k$\$\beta$ dimension $p$ $k$ dc $k$ $k$
border 24 \ \ 7 840 13 1.0 29 2
mfea-fac 6 \ \ 8 1000 9 1.4 45 18
Kdd9 21 \ \ 2 1000 14 1.0 NA 100
landsatEW 43 \ \ 6 1000 5 1.4 NA 99
balance-scale 44 \ \ 4 625 11 1.7 NA 84
pengleukEW 3 \ \ 10 72 11 1.0 8 43
Pendigits 23 \ \ 7 1000 24 1.0 105 91
energy-y2 20 \ \ 8 766 9 1.0 28 15
optical_test 12 \ \ 9 1000 10 1.9 42 3
soybean_test 4 \ \ 6 376 49 1.5 19 25
car 8 \ \ 6 1000 31 1.1 42 31
semeionEW 3 \ \ 6 1000 5 1.0 40 5
vote 46 \ \ 6 435 33 2.0 NA 97
Table 3: Hyperparameter values of the results on real-world datasets.
@@
Table 2 shows the results (bold means best) of our algorithm and the other eight comparison algorithms on real-world datasets. Table 3 shows their corresponding hyperparameter values. The ”\” indicates that the corresponding algorithm does not require hyperparameters, while ”NA” indicates that the algorithm fails to produce results. MDMSC achieves the best performance on most datasets. For example, on the high-dimensional dataset pengleukEW, MDMSC significantly outperforms most comparison algorithms. We attribute this to the consideration of manifold curvature, allowing it to effectively handle complex high-dimensional datasets. Additionally, on the large-scale dataset Pendigits, our algorithm also shows superior performance, demonstrating its capability to effectively handle large-scale datasets by partitioning pseudo-clusters. Although our algorithm does not always achieve the best results on a few datasets like landsatEW and soybean_test, its performance is very close to the optimal ones. This indicates that our algorithm possesses strong adaptability and effectiveness across a wide range of application scenarios.
Besides, We first performed a Friedman test on the ACC matric of MDMSC and the comparison algorithms. After confirming significant differences between the algorithms, we conducted a Nemenyi post-hoc test to compute the statistical significance of the differences between MDMSC and the comparison algorithms. We obtained $FriedmanStatistic=49.37$ with $P=2.10e-08<0.05$, confirming significant differences between the algorithms. In the Nemenyi post-hoc test, the critical difference threshold $CD=2.10$. Fig. 5 illustrates the mean rank of all algorithms, where a rank difference greater than CD indicates a significant difference between the two algorithms.
Figure 5: Mean rank of all algorithms on the ACC metric.